A 2 kilogram mass accelerates for 5 seconds persecond squared for 3 seconds. What is the impluse on the mass?
1
Expert's answer
2018-01-31T05:20:07-0500
J = F × t = m × a × t, where m – mass, a – acceleration, t – time;
J = 2×5×3 = 30 (N×s)